# Preparation of Papers for IEEE Sponsored Conferences & Symposia*

### v1

@@ V-A Soundness
Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ and intersection interval $\bar{I^{c}}$. For some $\delta<|\bar{I^{c}}|$, agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for $t\in[t^{i},t^{i}+\delta)$. The equivalent constraint set of forbidding constraints is $C_{i}=\left\{\langle i,m^{i},t\rangle\mid t\in[t^{i},t^{i}+\delta)\right\}$. As this is a shifting constraint, we know that a collision occurs with $j$ located at $v(w_{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. On the other hand, agent $j$ is forbidden from executing any wait action that requires occupying $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, or any move action $m$ with $\mathit{from}(m)=v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. All of these forbidden timed actions require being located at $v(w^{j})$ at some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. Thus, a collision will occur if $i$ and $j$ each execute any of their forbidden actions. Let $C_{j}$ contain one forbidding constraint for each of $j$’s forbidden actions. Then, it holds for any $\langle c_{i},c_{j}\rangle\in\langle C_{i},C_{j}\rangle$ that a valid solution must satisfy at least $c_{i}$ or $c_{j}$. Therefore, $\langle C_{i},C_{j}\rangle$ is sound and by Lemma III.1 so is our branching rule. ∎
@@
To conclude, our branching rule is shown to be sound when applied to move-move conflicts (in [20]) and move-wait conflicts (in Theorem V.1). Therefore, the branching rule is sound.
V-B A Requirement for Non-termination
@@
To prove that CCBS under our branching rule is solution complete and therefore terminates on a solvable MAPFR problem, we employ a proof by contradiction. We begin by assuming the opposite: that CCBS never terminates on some solvable MAPFR problem, despite the existence of a solution, and arrive at a condition which must hold if CCBS never terminates.
@@
We will establish in this section that if CCBS runs indefinitely, then the CT must contain an infinite sequence of nodes that are continuously selected and branched upon. By analyzing this infinite sequence, we derive a specific requirement that any non-terminating execution must satisfy. It will later be show in Section V-C that CCBS under our branching rule does not satisfy this requirement, thus proving termination. Specifically, what we establish here is:
@@
  1. V-B1) Infinite CT node sequence: There exists an infinite descending path in the constraint tree where each node contains conflicts and has objective value $\leq\sigma^{*}$, where $\sigma^{*}$ is the optimal.
@@
  2. V-B2) Concentration on $\tau^{\infty}$: Infinitely many nodes in the sequence map to the same trajectory $\tau^{\infty}$.
@@
  3. V-B3) Unbounded Constraint Accumulation: Constraints which determine how nodes map to $\tau^{\infty}$ accumulate unboundedly along the infinite node sequence, yet there must always exist another node mapping to $\tau^{\infty}$ and still maintaining an objective value lower than the optimal value.
@@
The final point formulates the requirement on the branching rule for non-termination.
We begin by defining trajectories and how plans map to them.
Definition V.2 (Trajectory).
@@
Definition V.3 (Joint Trajectory).
@@
Definition V.4 (Plan-Trajectory Mapping).
A plan $\pi$ maps to a trajectory $\tau$, denoted $\pi\sim\tau$, if $\tau$ contains exactly every move action in $\pi$ in the order that they appear. Similarly, $\Pi\sim\mathcal{T}$ if it holds for every agent $i$ that $\pi_{i}\sim\tau_{i}$ with $\pi_{i}\in\Pi,\tau_{i}\in\mathcal{T}$.
Informally, a trajectory $\tau$ describes the sequence of moves an agent takes when following a plan $\pi\sim\tau$, but without timing details. Multiple plans can therefore map to the same trajectory, as plans can contain various wait actions. For example, the following two plans
  • $\pi_{i}=\langle\langle m_{1}^{i},0\rangle,\langle w_{2}^{i},1.5\rangle,\langle m_{3}^{i},2\rangle,\langle m_{4}^{i},3\rangle,\langle w_{5}^{i},4.3\rangle\rangle$ from Figure 2, and
@@
both map to the same trajectory $\tau_{i}=\langle m_{1}^{i},m_{3}^{i},m_{4}^{i}\rangle$. Since each trajectory $\tau$ has a duration $\tau_{D}$, a joint trajectory $\mathcal{T}$ has an objective value $\sigma(\mathcal{T})$.
@@
Finally, we observe that the set $\mathbb{T}$ of all possible joint trajectories $\mathcal{T}$ is countable. This follows from the trajectories $\tau$ being finite sequences from the countable set $\mathcal{A}$, meaning that the set of all $\tau$ is countable. Since the set of all $\tau$ is countable, and a joint trajectory $\mathcal{T}$ comprises a discrete number of trajectories (exactly one per agent), the set of all possible joint trajectories $\mathbb{T}$ is countable. We define the subset $\mathbb{T}^{c}\subset\mathbb{T}$ to contain all joint trajectories with an objective value at or below some fixed constant $c\in\Re$:
@@
V-B1 Infinite CT node sequence
Since our branching rule is sound (Section V-A), all solutions (including the optimal solution) is reachable from the CT root. Non-termination implies that at every iteration, a non-solution CT node $N$ is expanded instead of a node representing an optimal solution with objective value $\sigma^{*}$. Since the search is best-first, it must hold that $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$ otherwise it would not be selected for expansion. Additionally, $N_{\Pi}$ must contain a conflict, otherwise $N_{\Pi}$ would constitute a valid solution. For non-termination to persist indefinitely, such nodes must be available at every iteration. This leads to the following recursive requirement: at every iteration there must exist at least one node $N$ satisfying the above two conditions, with a child node also satisfying the above two conditions, otherwise eventually all nodes satisfying these properties will be depleted and the algorithm will terminate with an optimal solution. In summary, non-termination implies that at every iteration exists at least one CT node $N$ where
@@ V-B1 Infinite CT node sequence
    $N_{\Pi}$ contains a conflict, and
@@
    a child node of NN satisfies (I) and (II).
@@
V-B2 Concentration on $\tau^{\infty}$
Consider a node $N^{k}\in\varsigma$. When $N^{k}$ is expanded, CCBS detects a conflict between two plans in $N^{k}_{\Pi}$ and applies the branching rule to create two child nodes. By definition of sequence $\varsigma$, exactly one of these children continues the sequence as $N^{k+1}$. For one of the plans involved in the conflict, say $\pi_{i}\in N^{k}_{\Pi}$, a constraint is created for agent $i$ and added to the constraint set of $N^{k+1}$. In the context of our analysis, we say that $\pi_{i}$ is “branched on” at $N^{k}$.
Definition V.5 (Node-trajectory mapping).
@@ V-B2 Concentration on $\tau^{\infty}$
With Definition V.5, we can map every one of the infinitely many nodes in $\varsigma$ to a specific trajectory. We now establish that infinitely many nodes in $\varsigma$ map to a specific trajectory, denoted $\tau^{\infty}$. Consider $N\in\varsigma$. The set of plans $N_{\Pi}$ satisfies $N_{\Pi}\sim\mathcal{T}$ for some $\mathcal{T}\in\mathbb{T}$, and $N$ maps to a trajectory $\tau\in\mathcal{T}$. By inequality (8), $\sigma\left(\mathcal{T}\right)\leq\sigma\left(N_{\Pi}\right)$, and by condition (I), $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$. Therefore, it follows that $\sigma\left(\mathcal{T}\right)\leq\sigma^{*}$, implying $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$. Since $\mathbb{T}^{\sigma^{*}}$ is finite and each $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$ contains a finite number of trajectories, the set of all possible trajectories that nodes in $\varsigma$ map to is finite. However, $\varsigma$ contains infinitely many nodes. By the Pigeonhole principle, infinitely many nodes in $\varsigma$ must map to at least one specific trajectory $\tau^{\infty}$.
V-B3 Unbounded Constraint Accumulation
For every branching at a node $N^{k}\in\varsigma$, a constraint is created and propagated to all subsequent nodes $N^{h}\in\varsigma$ for $h>k$. Since infinitely many nodes in $\varsigma$ map to $\tau^{\infty}$, constraints from branching on plans $\pi\sim\tau^{\infty}$ accumulate unboundedly, yet there must always exist some $\pi^{\prime}\sim\tau^{\infty}$ permitted under these constraints. On the other hand, by condition (I), every $N^{k}\in\varsigma$ must satisfy $\sigma\left(N^{k}_{\Pi}\right)\leq\sigma^{*}$. Since $\sigma$ is strictly monotonically increasing with the maximum plan duration, there exists a fixed constant $c<\infty$ (determined by $\sigma^{*}$) such that all plans in $N^{k}_{\Pi}$ have duration $\leq c$. Therefore, it must hold for all $N^{k}\in\varsigma$ that $\forall\pi\in N^{k}_{\Pi}:\pi_{D}\leq c$, otherwise $\sigma\left(N^{k}_{\Pi}\right)$ grows unboundedly to eventually violate condition (I).
Finally, we have arrived at our requirement for non-termination: for some fixed $c<\infty$ and trajectory $\tau^{\infty}$, infinitely many CCBS branching steps on a plan $\pi\sim\tau^{\infty}$ must preserve the existence of at least one valid plan $\pi^{\prime}\sim\tau^{\infty}$ with duration $\pi_{D}\leq c$.
V-C Solution Completeness
In this section, we prove that CCBS under our branching rule terminates within finite iterations by showing that the non-termination requirement established in Section V-B cannot be satisfied. This completes our proof by contradiction.
We prove that CCBS under our branching rule does not satisfy the non-termination requirement by showing that each branching operation reduces the set of feasible execution times for actions in $\tau^{\infty}$ by non-degenerate amounts, eventually making it impossible to construct any plan $\pi\sim\tau^{\infty}$ with a bounded duration.
@@ V-C Solution Completeness
  1. V-C1) Permitted Execution Times: We formalize how constraints restrict move execution and vertex occupancy times, establishing the framework for analyzing constraint accumulation.
  2. V-C2) CSIPP Planning Behavior: We explain how CCBS’s underlying path planner (CSIPP) constructs plans, focusing on how it selects execution times from feasible intervals.
  3. V-C3) Non-degenerate Reductions For Move Actions: We show how our branching rule applied to a timed move action reduces feasible move execution times by a non-degenerate amount.
  4. V-C4) Non-degenerate Reductions For Wait Actions: We show how our branching rule applied to a timed wait action reduces feasible move execution or feasible vertex occupancy times by a non-degenerate amount.
  5. V-C5) Bounded Plan Impossibility: Using the results from Sections V-C3 and V-C4, and the fact that all timed actions in a bounded-duration plan must end before a fixed time, we show that the feasible move execution and vertex occupancy times are exhausted after finite iterations of the CCBS search. Therefore, we show that CCBS under our branching rule does not satisfy the non-termination requirement and therefore is guaranteed to terminate.
V-C1 Permitted Execution Times
To prove that the non-termination requirement cannot be satisfied, we establish a framework to analyze how constraints accumulate and restrict agent movements. At the CT root $N^{R}$ where the constraint set is empty, $N^{R}_{C}=\varnothing$, all agents are permitted to execute all actions at any time $t\geq 0$. However, at nodes further down in the CT with non-empty constraint sets, any number of constraints may restrict when agents can execute moves and occupy vertices. Thus, under the constraints $N_{C}$ at CT node $N$, we define for a specific agent $i$

### v2

@@ 5.1 Soundness
Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}$, on which we apply $\delta$-BR. For agent $i$, $\delta$-BR creates the motion constraint $\langle i,m^{i},[t^{i},t^{i}+\delta)\rangle$. Since the motion constraint is in the form of a shifting constraint, and by (7) we know that $\delta<|\bar{I^{c}}|$, we know that a collision occurs if agent $j$ occupies vertex $v(w^{j})$ at any time during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. The motion constraint $\langle i,m^{i},[t^{i},t^{i}+\delta)\rangle$ can equivalently be expressed in its decomposed form — a set of forbidding constraints — $C_{i}=\left\{\langle i,m^{i},t\rangle\mid t\in[t^{i},t^{i}+\delta)\right\}$.
On the other hand, $\delta$-BR creates for agent $j$ a vertex constraint $\langle j,v(w^{j}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and the motion constraints $\langle j,m,[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ for every $m\in\mathcal{A}:\mathit{from}(m)=v(w^{j})$. These can collectively be represented by the set of forbidding constraints
$C_{j}$ $=\left\{\langle j,w,t\rangle\mid v(w)=v(w^{j}),[t,t+w_{D}]\cap[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\neq\varnothing\right\}$
$\cup\left\{\langle j,m,t\rangle\mid\mathit{from}(m)=v(w^{j}),m\in\mathcal{A},t\in[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\right\}$
For every constraint $\langle j,a,t\rangle\in C_{j}$, the forbidden timed action $\langle a,t\rangle$ requires $j$ being located at $v(w^{j})$ at some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$; the forbidden timed wait actions require $j$ occupying $v(w^{j})$ during some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ and the forbidden timed move actions start at $v(w^{j})$ at some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. Thus, if agent $i$ executes any of the actions forbidden by $C_{i}$ and agent $j$ executes any of the actions forbidden by $C_{j}$, then a collision will occur.
The above implies that for every $\langle c_{i},c_{j}\rangle=\langle\langle i,a^{i},t^{i}\rangle,\langle j,a^{j},t^{j}\rangle\rangle\in C_{i}\times C_{j}$ the pair of timed actions $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle$ is a conflict. Thus, any solution must satisfy $c_{i}\vee c_{j}$, meaning that $\langle c_{i},c_{j}\rangle$ is sound for every $\langle c_{i},c_{j}\rangle\in C_{i}\times C_{j}$ and therefore that $\langle C_{i},C_{j}\rangle$ is sound. By Lemma 3.1, this means that $\delta$-BR is sound on any move-wait conflict. ∎
@@
To conclude, $\delta$-BR is shown to be sound when applied to move-move conflicts (in [21]) and move-wait conflicts (in Theorem 5.1). As these are the only two conflict types considered, $\delta$-BR is therefore sound.
5.2 A Requirement for Non-termination
@@
To prove that CCBS using $\delta$-BR is solution complete and therefore terminates on a solvable MAPFR problem, we employ a proof by contradiction. We begin by assuming the opposite: that CCBS never terminates on a solvable MAPFR problem despite the existence of a solution. This will result in a condition which must hold if CCBS never terminates.
@@
We will establish in this section that if CCBS runs indefinitely, then the CT must contain an infinite sequence of nodes that are continuously selected and branched upon. By analyzing this infinite sequence, we derive a specific requirement that any non-terminating execution must satisfy. It will later be show in Section 5.3 that CCBS using $\delta$-BR does not satisfy this requirement, thus proving termination. Specifically, what we establish here is:
@@
  1. 5.2.1) Infinite CT node sequence: There exists an infinite descending path in the constraint tree where each node contains conflicts and has objective value $\leq\sigma^{*}$, with $\sigma^{*}$ being the optimal.
@@
  2. 5.2.2) Concentration on $\tau^{\infty}$: We define a trajectory in Definition 5.2 and a mapping from a CT node to a trajectory in Definition 5.5, and find that infinitely many nodes in the sequence map to the same trajectory $\tau^{\infty}$.
@@
  3. 5.2.3) Unbounded Constraint Accumulation: Constraints which determine how nodes map to $\tau^{\infty}$ accumulate unboundedly along the infinite node sequence, yet there must always exist another node mapping to $\tau^{\infty}$ which still maintains an objective value lower than the optimal value.
@@
The final point formulates the requirement on the branching rule for non-termination. We begin by defining trajectories and how plans map to them.
Definition 5.2 (Trajectory).
@@
Definition 5.3 (Joint Trajectory).
@@
Definition 5.4 (Plan-Trajectory Mapping).
A plan $\pi$ maps to a trajectory $\tau$, denoted $\pi\sim\tau$, if $\tau$ contains exactly every move action in $\pi$ in the order that they appear. Similarly, a joint plan $\Pi$ maps to a joint trajectory $\mathcal{T}$, $\Pi\sim\mathcal{T}$, if it holds for every agent $i$ that $\pi_{i}\sim\tau_{i}$ with $\pi_{i}\in\Pi,\tau_{i}\in\mathcal{T}$.
Informally, a trajectory $\tau$ describes the sequence of move actions an agent takes when following a plan $\pi\sim\tau$, but without timing details. Multiple plans can therefore map to the same trajectory, as plans can contain various wait actions. For example, the following two plans
  • $\pi_{i}=\langle\langle m_{1}^{i},0\rangle,\langle w_{2}^{i},1.5\rangle,\langle m_{3}^{i},2\rangle,\langle m_{4}^{i},3\rangle,\langle w_{5}^{i},4.3\rangle\rangle$ introduced in Figure 2, and
@@
both map to the same trajectory $\tau_{i}=\langle m_{1}^{i},m_{3}^{i},m_{4}^{i}\rangle$. Since each trajectory $\tau$ has a duration $\tau_{D}$, a joint trajectory $\mathcal{T}$ can be evaluated on the objective function, $\sigma(\mathcal{T})$.
@@
Finally, we observe that the set $\mathbb{T}$ of all possible joint trajectories $\mathcal{T}$ is countable. This follows from the trajectories $\tau$ being finite sequences from the countable set $\mathcal{A}$, meaning that the set of all $\tau$ is countable. Since the set of all $\tau$ is countable, and a joint trajectory $\mathcal{T}$ comprises a discrete number of trajectories (exactly one per agent), the set $\mathbb{T}$ is countable. We define the subset $\mathbb{T}^{c}\subset\mathbb{T}$ to contain all joint trajectories with an objective value at or below some fixed constant $c\in\Re$:
@@
5.2.1 Infinite CT node sequence
Since $\delta$-BR is sound (Section 5.1), all solutions (including all optimal solutions) are reachable from the CT root. Non-termination of CCBS implies that at every iteration, a non-solution CT node $N$ is expanded instead of a node representing an optimal solution with objective value $\sigma^{*}$. Since the search is best-first, it must hold that $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$ otherwise it would not be selected for expansion. Additionally, $N_{\Pi}$ must contain a conflict, otherwise $N_{\Pi}$ would constitute a valid solution. For non-termination to persist indefinitely, such nodes must be available at every iteration. This leads to the following recursive requirement: at every iteration there must exist at least one node $N$ satisfying the above two conditions, with at least one child node also satisfying the above two conditions, otherwise eventually all nodes satisfying these properties will be depleted and the algorithm will terminate with an optimal solution. In summary, non-termination implies that at every iteration there exists at least one CT node $N$ where
@@ 5.2.1 Infinite CT node sequence
    $N_{\Pi}$ contains conflicts and is therefore not a solution, and
@@
    at least one child node of NN satisfies (I) and (II).
@@
5.2.2 Concentration on $\tau^{\infty}$
Consider a node $N^{k}\in\varsigma$. When $N^{k}$ is expanded, CCBS detects a conflict between two plans in $N^{k}_{\Pi}$ and applies a branching rule to create two child nodes. By definition of sequence $\varsigma$, exactly one of these children continues the sequence as $N^{k+1}$. For one of the plans involved in the conflict, say $\pi_{i}\in N^{k}_{\Pi}$, a constraint is created for agent $i$ and added to the constraint set $N^{k+1}_{C}$. In the context of our analysis, we say that $\pi_{i}$ is “branched on” at $N^{k}$.
Definition 5.5 (Node-trajectory mapping).
@@ 5.2.2 Concentration on $\tau^{\infty}$
With Definition 5.5, we can map every one of the infinitely many nodes in $\varsigma$ to a specific trajectory. We now establish that infinitely many nodes in $\varsigma$ map to a specific trajectory, denoted $\tau^{\infty}$. Consider $N\in\varsigma$. The joint plan $N_{\Pi}$ satisfies $N_{\Pi}\sim\mathcal{T}$ for some $\mathcal{T}\in\mathbb{T}$, and $N$ maps to a trajectory $\tau\in\mathcal{T}$. By inequality (8), $\sigma\left(\mathcal{T}\right)\leq\sigma\left(N_{\Pi}\right)$, and by condition (I), $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$. Therefore, it follows that $\sigma\left(\mathcal{T}\right)\leq\sigma^{*}$ and therefore that $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$. Since $\mathbb{T}^{\sigma^{*}}$ is finite and each $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$ contains a finite number of trajectories, the set of all possible trajectories that nodes in $\varsigma$ map to is finite. However, $\varsigma$ contains infinitely many nodes. By the Pigeonhole principle, there exists at least one trajectory $\tau^{\infty}$ which infinitely many nodes in $\varsigma$ map to.
5.2.3 Unbounded Constraint Accumulation
For every branching at a node $N^{k}\in\varsigma$, a constraint $c$ is created and propagated to all subsequent nodes $N^{h}\in\varsigma$ for $h>k$, such that $c\in N^{h}_{C}$. Since infinitely many nodes in $\varsigma$ map to $\tau^{\infty}$, constraints from branching on plans $\pi\sim\tau^{\infty}$ accumulate unboundedly along $\varsigma$, yet there must always exist some $\pi^{\prime}\sim\tau^{\infty}$ permitted under these constraints. On the other hand, by condition (I), every $N^{k}\in\varsigma$ must satisfy $\sigma\left(N^{k}_{\Pi}\right)\leq\sigma^{*}$. Since $\sigma$ is strictly monotonically increasing with the maximum plan duration, there exists a fixed constant $c<\infty$ (determined by $\sigma^{*}$) such that all plans in $N^{k}_{\Pi}$ have duration $\leq c$. Therefore, it must hold for all $N^{k}\in\varsigma$ that $\forall\pi\in N^{k}_{\Pi}:\pi_{D}\leq c$, otherwise $\sigma\left(N^{k}_{\Pi}\right)$ necessarily grows to eventually violate condition (I).
Finally, we have arrived at our requirement for non-termination: for some fixed $c<\infty$ and trajectory $\tau^{\infty}$, there must exists some plan $\pi\sim\tau^{\infty}$ with duration $\pi_{D}\leq c$ that is permitted under the constraints accumulated from infinitely many branchings on plans also mapping to $\tau^{\infty}$.
5.3 Solution Completeness
In this section, we prove that CCBS under $\delta$-BR terminates within finite iterations by showing that the non-termination requirement established in Section 5.2 cannot be satisfied, thereby completing our proof by contradiction. This is done by showing that each branching operation reduces the set of feasible execution times for actions in $\tau^{\infty}$ by non-degenerate amounts, eventually making it impossible to construct any plan $\pi\sim\tau^{\infty}$ with a bounded duration.
@@ 5.3 Solution Completeness
  1. 5.3.1) Permitted Execution Times: We formalize how constraints produced under $\delta$-BR restrict move execution and vertex occupancy times, establishing the framework for analyzing constraint accumulation.
  2. 5.3.2) CSIPP Planning Behavior: We explain how CCBS’s underlying path planner, CSIPP, constructs plans, focusing on how it selects execution times from safe intervals.
  3. 5.3.3) Non-degenerate Reductions For Move Actions: We show how $\delta$-BR applied to a timed move action (in a move-move or move-wait conflict) reduces feasible move execution times by a non-degenerate amount.
  4. 5.3.4) Non-degenerate Reductions For Wait Actions: We show how $\delta$-BR applied to a timed wait action reduces feasible move execution or feasible vertex occupancy times by a non-degenerate amount.
  5. 5.3.5) Bounded Plan Impossibility: Using the results from Sections 5.3.3 and 5.3.4, and the fact that all timed actions in a bounded-duration plan must end before a fixed time, we show that the feasible move execution and vertex occupancy times are exhausted after finite iterations of the CCBS search. Therefore, we show that CCBS under $\delta$-BR does not satisfy the non-termination requirement and therefore is guaranteed to terminate.
5.3.1 Permitted Execution Times
To prove that the non-termination requirement cannot be satisfied, we establish a framework to analyze how constraints accumulate and restrict agent movements. At the CT root $N^{R}$ where the constraint set is empty, $N^{R}_{C}=\varnothing$, all agents are permitted to execute all actions at any time $t\geq 0$. However, at nodes further down in the CT with non-empty constraint sets, any number of constraints may restrict when agents can execute move actions and occupy vertices. Thus, under the constraints $N_{C}$ at CT node $N$, we define for a specific agent $i$